Let’s consinder a mortgage application using HMDA (The Home Mortgage Disclosure Act). Here is a sample from 30 mortgage applications.

ID

loanamt

income

hprice

1

109

63

155

2

185

137

264

3

121

53

128

4

125

78

125

5

119

37

149

6

153

65

171

7

380

188

484

8

100

58

125

9

110

78

158

10

41

31

116.5

11

115

54

128

12

248

117

280

13

126

60

157.5

14

260

192

325

15

90

40

145

16

50

36

230

17

125

45

125

18

125

55

145

19

158

62

175

20

130

29

209

21

204

77

260

22

30

28

150

23

114

60

143

24

188

91

253

25

187

85

285

26

84

44

105

27

450

265

650

28

108

49

120

29

100

53

125

30

53

24

66

loanamt = Amount of Mortgage Loan Application (in $1000)

income = Applicant’s Annual Income (in $1000)

hprice = House Price to buy (in $1000)

Regression Analyis

Let’s consider the following regression model. Estimate the model using Minitab and answer the questions using the output.

Loanamti = b0 + b1 * incomei + et

6) t test for the coefficient of income (Ho: B1 = 0 )

7) F statistics and perform the test for the model

8) Variance of et

9) According to the model what are the predicted loan amount if applicants have annual income of $50,000, $100,000, and $200,000 and their confidence intervals?

10) List and explain the assumptions you made for a simple regression model
Answer #1

As asked, using minitab'17 the analysis would be as followed:

Using the above model as; loanamt= b0+ b1 income

6)

In minitab, computing of the linear regression model also gives the t-value for both the constant as well as coefficient.

Here, hypothesis would be as followed:

H0: b0=0

H1: b0≠0

The analysis from minitab is :

Session Coefficients VIF Term Constant income Coef SE Coef T-Value 29.4 10.5 2.81 1.556 0.113 13.74 P-value 0.009 0.000 1.00t-value corresponding to income is 13.74 and p-value corresponding to it is <0.0001 , which means we would reject the null hypothesis.

7) By using ANOVA table, the computed F-statistic would be as followed:

Hypothesis would be as followed:

Method Null hypothesis All means are equal Alternative hypothesis At least one mean is different significance level a = 0.05

il Minitab - Untitled file Edit Data Calc Stat Graph Editor Tools Window Help Assistant #H% 11 03 SO GOODSOD GIF | -- A3 KG +

As the p-value is too low, so we may reject the null hypothesis.

8) Variance of et refers to the variance of the residuals; residuals calculated using the minitab would be as followed:

Worksheet 1 *** C8 C3 hprice C6 Variance1 1077.96 C7 N1 30 137 78 37 C1 C2 loanamt income 109 63 185 121 125 119 153 65 380 1

It's variance is 1077.96

9)

predicted loan amount would be calculated by using the following equation:

Model Summary S R -sq 33.4135 87.09% R-sq(adj) 86.63% R-sq (pred) 84.308 Coefficients VIF Term Constant income Coef SE Coef 2

Put the values in the place of income you'll get your predicted amount from the equation as followed:

loanamt= 29.4+1.556*50000= 77829.4

loanamt= 29.4+1.556*100000= 155629.4

loanamt= 29.4+1.556*200000= 311229.4